OpenStudy (anonymous):

evaluate the integral: int: [(x^3 + 3x^2 + x + 9)/(x^2 + 1)(x^2 + 3) ] dx using the linear quotient rule I have: let x^3 + 3x^2 + x + 9 = Ax+B (x^2 + 3) + Cx + D(x^2 + 1) I am trying to solve for A B C and D...

OpenStudy (anonymous):

You have to us the method of partial fractions: \[ \frac{x^3+3 x^2+x+9}{\left(x^2+1\right) \left(x^2+3\right)}=\frac{A x+B}{x^2+1}+\frac{C x+D}{x^2+3} \] After finding A, B ,C, D, the integral would become easy.

OpenStudy (anonymous):

Use the cover-up method.

OpenStudy (anonymous):

You will have to the work. To check your answer, A=0; B=3; C=1 and D=0

OpenStudy (anonymous):

Your integral is now \[ \int \frac{x^3+3 x^2+x+9}{\left(x^2+1\right) \left(x^2+3\right)} \, dx=\int \frac{3}{x^2+1} \, dx+\int \frac{x}{x^2+3} \, dx=\\ 3 \tan^{-1} x + \frac 1 2 \ln(x^2 +3) + K \]

OpenStudy (anonymous):

I got it. Thank you so much for your help. A= 0 B = 3 C = 1 D = 0.

OpenStudy (anonymous):

yw